Grade 11Mechanics

A particle experiences constant acceleration for 6 seconds after starting from rest if it travels a distance d1 in first 2 seconds D2 in next 2 seconds and a distance d 3 in last 2 seconds then what is the ratio of the D1 is to D2 to D3

rubiks cuber
8 Years agoGrade 11

1 Answer

Nain aditya
8 Years ago
As the particle is satrting from rest the initial velocity = 0We know that s = ut+1/2 at^2So , d1 is distance travelled after 2 second D1= 0*2+1/2a2^2I.e, D1= 2aAnd D2 is distance travelled in next 2 second so D2 = distance travelled in 4 sec - D1I.e, D2 = (0*4 + 1/2a4^2)-2aD2= 8a-2a= 6aD3 is distance travelled by the particle on last 2 sec so , D3 = distance travelled in 6 second - distance travelled in 4 secI.e D3 =( 0*6+1/2a6^2 ) - 8aD3 = 18a-8a = 10a So D1:D2:D3 = 2a:6a:10a=1:3:5